find two consecutive integers such that the sum of 7 times the first integer and 4 times the second integer is 37
Answer by checkley71(8403)

7x+4(x+1)=37
7x+4x+4=37
11x=37-4
11x=33
x=33/11
x=3 for the smaller number
3+1=4 the second number.
proof
7*3+4*4=37
21+16=37
37=37